OpenStudy (anonymous):

Solve for the roots in the following equation. Hint: Factor both quadratic expressions. (x^4 + 5x^2 - 36)(2x^2 + 9x - 5) = 0

OpenStudy (janu16):

ok so Which two numbers add up to 5 and multiply to -36?

OpenStudy (anonymous):

in the equation?

OpenStudy (janu16):

yes for the first part

OpenStudy (janu16):

ok so what is -4 plus 9?

OpenStudy (anonymous):

five =)

OpenStudy (janu16):

and does those two numbers multiplies to -36?

OpenStudy (anonymous):

indeed

OpenStudy (janu16):

so when we rewrite its gonna be : (x^2-4)(x^2+9)(2x^2 + 9x - 5) = 0

OpenStudy (anonymous):

okay

OpenStudy (janu16):

now x^2-4 fits the form a^2 - b^2 so we are gonna write it like this (x^2-2^2)(x^2+9)(2x^2 + 9x - 5) = 0

OpenStudy (anonymous):

oki

OpenStudy (janu16):

we are gonna use difference of square so a^2-b^2=(a+b)(a-b)

OpenStudy (janu16):

so (x+2)(x-2)(x^2+9)(2x^2 + 9x - 5) = 0

OpenStudy (janu16):

Multiply 2 by -5, which is -10. Which two numbers add up to 9 and multiply to -10?

OpenStudy (janu16):

what do you think?

OpenStudy (janu16):

dan

OpenStudy (anonymous):

sorry ill brb my father needs help

OpenStudy (janu16):

k

OpenStudy (anonymous):

hello im back

OpenStudy (anonymous):

in the first and the last

OpenStudy (welshfella):

so we are factoring 2x^2 + 9x - 5 right?

OpenStudy (anonymous):

um maybe i really dont get any of this stuff

OpenStudy (welshfella):

ok these trinomials factor to the form (ax + b(bcx + d) whete a b c and d are integers now we have 2x^2 + 9x - 5 so start we must have 2x in one bracket and x in the other (2x + b)(x + d) because the 2x 8 x will give us the 2x^2 do you follow that ok?

OpenStudy (welshfella):

* 2x * x = 2x^2